Contributed talks

Ahn, Eunkyung: Thursday, June 19, 17:45, Room 4 Al Zhour, Zeyad: Thursday, June 19, 12:15, Room 4 Andjelic, Milica: Friday, June 20, 10:35, Room 3 Arav, Marina: Tuesday, June 17, 10:35, Room 3 Armandnejad, Ali: Tuesday, June 17, 12:15, Room 3 Barría, José: Friday, June 20, 16:20, Room 4 Bourgeois, Gerald: Monday, June 16, 17:20, Room 4 Bru, Rafael: Monday, June 16, 18:10, Room 3 Bueno Cachadina, María Isabel: Tuesday, June 17, 17:20, Room 4 Carriegos, Miguel: Tuesday, June 17, 11:50, Room 4 Castro-González, Nieves: Friday, June 20, 11:25, Room 4 Cheng, Wei: Friday, June 20, 16:45, Room 4 Cherepniov, Mikhail: Thursday, June 19, 11:50, Room 4 Corral, Cristina: Monday, June 16, 11:10, Room 3 Cortés, Vanesa: Friday, June 20, 15:30, Room 3 Costa, Liliana: Monday, June 16, 10:45, Room 4 Cravo, Glória: Tuesday, June 17, 10:35, Room 4 da Cruz, Henrique F.: Thursday, June 19, 18:35, Room 4 Day, Jane: Thursday, June 19, 16:55, Room 3 De Terán, Fernando: Monday, June 16, 17:45, Room 4 Djordjevic, Slavisa: Thursday, June 19, 11:00, Room 3 Dodig, Marija: Monday, June 16, 10:45, Room 3 Dolinar, Gregor: Thursday, June 19, 17:20, Room 4 Domínguez, María Elena: Monday, June 16, 17:20, Room 3 Dopazo, Esther: Friday, June 20, 10:35, Room 4 Fernandes, Rosário: Monday, June 16, 11:10, Room 4 Ferrer, Josep: Tuesday, June 17, 12:15, Room 4 Fonseca, Carlos: Thursday, June 19, 12:15, Room 3 Fosner, Ajda: Monday, June 16, 11:35, Room 3 Frank, Martin: Tuesday, June 17, 17:45, Room 4 Furtado, Susana: Thursday, June 19, 11:25, Room 3 Furuichi, Shigeru: Monday, June 16, 16:55, Room 4 Gassó, Maria T.: Friday, June 20, 11:00, Room 4 Glebsky, Lev: Tuesday, June 17, 11:50, Room 3 Goldberger, Assaf: Friday, June 20, 16:20, Room 3 Gouveia, María: Monday, June 16, 16:55, Room 3 Grudsky, Sergey: Monday, June 16, 18:35, Room 3 Guzmán, José Ramón: Thursday, June 19, 16:55, Room 5 Hanaish, Ibrahim: Thursday, June 19, 17:45, Room 5 Horn, Roger: Monday, June 16, 12:00, Room 3 Huylebrouck, Dirk: Monday, June 16, 18:10, Room 2 Im, Bokhee: Thursday, June 19, 10:35, Room 5 Kaibah, Hussein: Thursday, June 19, 18:10, Room 5 Kortesi, Peter: Monday, June 16, 17:20, Room 2 Lee, Gue Myung: Thursday, June 19, 17:20, Room 5 Lee, Hosoo: Monday, June 16, 18:35, Room 4 Machado, Silvia: Monday, June 16, 16:55, Room 2 Maroulas, John: Wednesday, June 18, 11:00, Room 4 Marovt, Janko: Thursday, June 19, 11:00, Room 4 Marques, Maria-da-Graça: Thursday, June 19, 16:55, Room 4 Martínez, José-Javier: Tuesday, June 17, 18:35, Room 4 Martins, Enide: Friday, June 20, 11:00, Room 3 Milligan, Thomas: Friday, June 20, 15:55, Room 4 Narayan, Sivaram: Friday, June 20, 12:15, Room 3 Olshevsky, Vadim: Thursday, June 19, 11:25, Room 4 Palma, Alejandro: Thursday, June 19, 11:00, Room 5 Patricio, Pedro: Friday, June 20, 11:50, Room 4 Peña, Marta: Tuesday, June 17, 11:00, Room 4 Poole, George: Monday, June 16, 17:45, Room 2 Prokip, Volodymyr: Thursday, June 19, 18:10, Room 4 Protasov, Vladimir: Wednesday, June 18, 11:50, Room 4 Pruneda, Rosa E.: Tuesday, June 17, 11:00, Room 3 Rakotondrajao, Fanja: Friday, June 20, 17:10, Room 4 Roca, Alicia: Tuesday, June 17, 11:25, Room 4 Rodríguez, Juan Alberto: Friday, June 20, 11:50, Room 3 Rump, Siegfried M.: Monday, June 16, 17:45, Room 3 Russo, Maria Rosaria: Tuesday, June 17, 16:55, Room 4 Salam, Ahmed: Tuesday, June 17, 18:10, Room 4 Satô, Kenzi: Thursday, June 19, 11:25, Room 5 Schaeffer, Elisa: Friday, June 20, 11:25, Room 3 Schaffrin, Burkhard: Friday, June 20, 12:15, Room 2 Sebeldin, Anatoly: Thursday, June 19, 11:50, Room 5 Seddighin, Morteza: Wednesday, June 18, 12:15, Room 4 Semrl, Peter: Tuesday, June 17, 11:25, Room 3 Sendov, Hristo: Thursday, June 19, 11:50, Room 3 Shahryari, Mohammad: Thursday, June 19, 12:15, Room 5 Sivakumar, Koratti Chengalrayan: Friday, June 20, 12:15, Room 4 Sivic, Klemen: Thursday, June 19, 10:35, Room 4 Soares, Graça: Friday, June 20, 15:30, Room 4 Stosic, Marko: Monday, June 16, 12:25, Room 3 Stuart, Jeffrey: Monday, June 16, 11:35, Room 4 Teixeira Matos, Isabel: Friday, June 20, 15:55, Room 3 Uchiyama, Mitsuru: Monday, June 16, 18:10, Room 4 Uhlig, Frank: Wednesday, June 18, 10:35, Room 4 Vallejo, Ernesto: Monday, June 16, 12:25, Room 4 Vargas, Xaab Nop: Monday, June 16, 18:35, Room 2 Weaver, James: Monday, June 16, 12:00, Room 4 Wojciechowski, Piotr: Friday, June 20, 16:45, Room 3 Wrobel, Iwona: Thursday, June 19, 10:35, Room 3 Wu, Pei Yuan: Wednesday, June 18, 11:25, Room 4